# Cooperative Nonlinear Guidance Strategies for Guaranteed Pursuit-Evasion

### v2

@@ III Main Results
$\ddot{\lambda}_{\mathrm{EP}}=$ $~{}\dfrac{\dot{\lambda}_{\mathrm{EP}}}{r_{\mathrm{EP}}}\left(-\dot{r}_{\mathrm{EP}}-v_{\mathrm{P}}\cos\delta_{\mathrm{PE}}+v_{\mathrm{E}}\cos\delta_{\mathrm{EP}}\right)$
$~{}+\dfrac{\cos\delta_{\mathrm{PE}}}{r_{\mathrm{EP}}}a_{\mathrm{P}}-\dfrac{\cos\delta_{\mathrm{EP}}}{r_{\mathrm{EP}}}a_{\mathrm{E}},$ (7)
which can be simplified using (2a) to
$\ddot{\lambda}_{\mathrm{EP}}=$ $~{}\dfrac{-2\dot{r}_{\mathrm{EP}}\dot{\lambda}_{\mathrm{EP}}}{r_{\mathrm{EP}}}+\dfrac{\cos\delta_{\mathrm{PE}}}{r_{\mathrm{EP}}}a_{\mathrm{P}}-\dfrac{\cos\delta_{\mathrm{EP}}}{r_{\mathrm{EP}}}a_{\mathrm{E}}.$ (8)
A similar procedure results in
$\ddot{\lambda}_{\mathrm{DP}}=$ $~{}\dfrac{-2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}a_{\mathrm{D}},$ (9)
$\ddot{\lambda}_{\mathrm{DE}}=$ $~{}\dfrac{-2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}+\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}-\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}a_{\mathrm{D}}.$ (10)
One may notice from (8)–(10) that the dynamics of LOS angles have a relative degree of two with respect to the lateral acceleration (steering controls) of the corresponding agents. This concludes the proof. ∎
Using the results in Lemma 1, we can now obtain the dynamics of the error variable, $\beta$, as
$\ddot{\beta}=$ $~{}\ddot{\lambda}_{\mathrm{DP}}-\ddot{\lambda}_{\mathrm{DE}}$ (11)
$=$ $~{}-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}$
$~{}+\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}},$ (12)
which is affine with the steering controls of the evader and the defender, whereas that of the pursuer is an unknown quantity. The dynamics of $\beta$ can be written as
$\ddot{\beta}=$ $~{}\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}+\mathcal{U},$ (13)
@@
$\mathcal{U}=-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}+\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}.$ (14)
In order to nullify $\beta$, we need to design $a_{\mathrm{D}}$ and $a_{\mathrm{E}}$ such that the evader-defender team cooperatively maneuvers to place the defender on a fixed $\chi^{\star}$ in the interval $\left[\dfrac{\pi}{2},\dfrac{3\pi}{2}\right]$ regardless of the initial three-body engagement geometry.
Toward this objective, we consider a dual-layer sliding manifold, with the inner-layer as
@@
for some $k_{2}\in\mathbb{N}$, and $h(0)$ is designed to ensure $\bar{\mathcal{S}}(0)=0$, which can be obtained from (17) as $h(0)=-\underaccent{\bar}{\mathcal{S}}(0)$. In (18), $t_{1}$ is the exact time instant from which onward the inner layer sliding manifolds become identically equal to zero. In other words, the sliding mode is enforced on the inner-layer sliding manifold at time $t=t_{1}$. The fundamental concept behind eliminating the reaching phase is to ensure global exact-time convergence for the sliding variables and achieve a complete sliding mode response. This, in consequence, bolsters robustness, as sliding mode controllers typically experience diminished robustness during their reaching phase. We are now ready to present the joint cooperative effort of the evader-defender alliance next.
Theorem 2.
Consider the three-body engagement described using (2), the dynamics of the error variable, (13), and the dual-layer sliding manifold in (15) and (17). If the joint cooperative maneuver of the evader-defender team is designed as
$\mathcal{U}=-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}-\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}+\dot{g}(t)-\dot{h}(t)-\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}}),$ (19)
@@
Consider a continuous, radially unbounded Lyapunov function candidate $V=|\bar{\mathcal{S}}|$. Hereafter, we streamline the notation again by dropping the arguments of variables denoting their time dependency. Upon differentiating $V$ with respect to time and using the relations in (15) and (17), one may obtain
$\dot{V}=\mathrm{sign}(\bar{\mathcal{S}})\dot{\bar{\mathcal{S}}}=\mathrm{sign}(\bar{\mathcal{S}})\left(\dot{\underaccent{\bar}{\mathcal{S}}}+\dot{h}\right)=\mathrm{sign}(\bar{\mathcal{S}})\left(\ddot{\beta}-\dot{g}+\dot{h}\right),$ (20)
which can be further written using the expression in (13) as
$\dot{V}=$ $\mathrm{sign}(\bar{\mathcal{S}})\left[-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{2\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{DE}}}a_{\mathrm{E}}\right.$
$~{}\left.+\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}-\dot{g}+\dot{h}\right].$ (21)
Upon substituting the expression of $\mathcal{U}$ from (19) in the above relation, the derivative of the Lyapunov function candidate reduces to
$\dot{V}=$ $\mathrm{sign}(\bar{\mathcal{S}})\left[-\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}})+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}\right]$
$\leq$ $-\mathcal{K}+\dfrac{1}{r_{\mathrm{DP}}}a_{\mathrm{P}}^{\max}~{}<0,$ (22)
@@
Note that $\mathcal{U}$ in Theorem 2 is the net control authority that the evader-defender pair has to maintain in order to achieve a successful interception of the pursuer. Next, we endeavor to allocate this net control effort to the evader and the defender. Needless to say, there can be several choices for the allocation of the total control effort between the defender and the evader. However, we focus on a control allocation technique that instantaneously minimizes the $\mathcal{L}_{2}$ norm of the net control subject to the affine constraint (14). Towards that, our aim is to minimize
$\mathcal{C}:=\sqrt{a_{\mathrm{E}}^{2}+a_{\mathrm{D}}^{2}}.$ (27)
The essence of this effort allocation is presented in the next theorem.
@@
The evader-defender team cooperatively ensures the interception of the pursuer with minimum values of the lateral accelerations,
$a_{\mathrm{E}}=$ $\dfrac{2r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}\left[r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)-r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}\right]}$
$+\dfrac{2r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}\left[r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)-r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}\right]}$
$-\dfrac{r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{g}(t)-r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{h}(t)-r_{\mathrm{DE}}r_{\mathrm{DP}}\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}})}{\left[r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)-r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}\right]},$ (28)
$a_{\mathrm{D}}=$ $\dfrac{2r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}\left[r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}-r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)\right]}$
$+\dfrac{2r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{r}_{\mathrm{DE}}\dot{\lambda}_{\mathrm{DE}}}{r_{\mathrm{DE}}\left[r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}-r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)\right]}$
$-\dfrac{r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{g}(t)-r_{\mathrm{DE}}r_{\mathrm{DP}}\dot{h}(t)-\mathcal{K}r_{\mathrm{DE}}r_{\mathrm{DP}}\mathrm{sign}(\bar{\mathcal{S}})}{\left[r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}-r_{\mathrm{DP}}\left(\cos\delta_{\mathrm{DE}}+\cos\delta_{\mathrm{ED}}\right)\right]}.$ (29)
@@
$a_{\mathrm{E}}=\dfrac{r_{\mathrm{DE}}}{\cos\delta_{\mathrm{ED}}}\left[\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{DE}}}-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}-\mathcal{U}\right].$ (30)
Now substituting the above expression of $a_{\mathrm{E}}$ in the cost function (27) yields which can also be written as $\mathcal{C}=\sqrt{\mathfrak{B}_{1}\left(\mathfrak{B}_{2}a_{\mathrm{D}}-\mathcal{U}\right)^{2}+a_{\mathrm{D}}^{2}}$, where $\mathfrak{B}_{1}=\dfrac{r_{\mathrm{DE}}}{\cos\delta_{\mathrm{ED}}}$ and $\mathfrak{B}_{2}=\dfrac{r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}-r_{\mathrm{DE}}\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DE}}r_{\mathrm{DP}}}$. Computing the partial derivative of $\mathcal{C}$ with respect to $a_{\mathrm{D}}$ and equating it to zero leads to
@@
$\dfrac{\partial\mathcal{C}}{\partial a_{\mathrm{D}}}=\dfrac{1}{2\mathcal{C}}\left[2\mathfrak{B}_{1}\left(\mathfrak{B}_{2}a_{\mathrm{D}}-\mathcal{U}\right)\mathfrak{B}_{2}+2a_{\mathrm{D}}\right]=0$
@@
$\implies\mathfrak{B}_{1}\mathfrak{B}_{2}^{2}a_{\mathrm{D}}+a_{\mathrm{D}}=\mathfrak{B}_{1}\mathcal{U}\implies a_{\mathrm{D}}=\dfrac{\mathfrak{B}_{1}\mathcal{U}}{1+\mathfrak{B}_{1}\mathfrak{B}_{2}^{2}},$ (31)

### v3

@@ 3 Main Results
$\ddot{\lambda}_{\mathrm{DP}}=$ $~\dfrac{\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}\left(-\dot{r}_{\mathrm{DP}}-v_{\mathrm{P}}\cos\delta_{\mathrm{PD}}+v_{\mathrm{D}}\cos\delta_{\mathrm{PD}}\right)+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}$
$~-\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}a_{\mathrm{D}},$ (8)
which can be simplified using (2c) to
$\ddot{\lambda}_{\mathrm{DP}}=$ $~\dfrac{-2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\left(\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{P}}-\left(\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}.$ (9)
By following a similar procedure, one may obtain $\ddot{\lambda}_{\mathrm{ED}}$ as
$\ddot{\lambda}_{\mathrm{ED}}=$ $~\dfrac{-2\dot{r}_{\mathrm{ED}}\dot{\lambda}_{\mathrm{ED}}}{r_{\mathrm{ED}}}-\left(\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{ED}}}\right)a_{\mathrm{E}}+\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{ED}}}\right)a_{\mathrm{D}}.$ (10)
Using (9) and (10), the expression in (5) becomes
$\ddot{\beta}$ $=\dfrac{2\dot{r}_{\mathrm{ED}}\dot{\lambda}_{\mathrm{ED}}}{r_{\mathrm{ED}}}-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\left(\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{ED}}}\right)a_{\mathrm{E}}+\left(\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{P}}$
$~-\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{ED}}}+\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}},$ (11)
which is affine with the steering controls of the evader and the defender, whereas that of the pursuer is an unknown quantity. ∎
We now consider two levels of cooperation (information-level and maneuver-level) between the evader and the defender, depending upon the communication between them. In information-level cooperation, the evader independently performs maneuvers to evade the pursuer, irrespective of the defender’s actions. However, the evader shares its maneuvering information with the defender. In other words, in information-level cooperation, the evader is indirectly cooperating with the defender when the former does not actively participate with the latter to attain the said angle $\chi^{\star}\in\left[{\pi}/{2},{3\pi}/{2}\right]$. In maneuver-level cooperation, the evader actively participates (maneuvers) with the defender to attain the said angle $\chi^{\star}\in\left[{\pi}/{2},{3\pi}/{2}\right]$. We first discuss the maneuver-level cooperation scenario between the evader and the defender, wherein the joint cooperative efforts of the evader-defender alliance are designed.
One can write the dynamics of $\beta$ given in (11) as
$\ddot{\beta}=$ $~\dfrac{2\dot{r}_{\mathrm{ED}}\dot{\lambda}_{\mathrm{ED}}}{r_{\mathrm{ED}}}-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}+\mathcal{U},$ (12)
@@
$\mathcal{U}=\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{ED}}}a_{\mathrm{E}}-\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{ED}}}+\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}.$ (13)
In order to nullify $\beta$, we need to design $a_{\mathrm{D}}$ and $a_{\mathrm{E}}$ such that the evader-defender team maneuvers to place the defender on a fixed $\chi^{\star}$ in the interval $\left[\pi/2,3\pi/2\right]$ regardless of the initial three-body engagement geometry.
Toward this objective, we consider a dual-layer sliding manifold, with the inner layer as
@@
for some $k_{2}\in\mathbb{N}$, and $h(0)$ is designed to ensure $\bar{\mathcal{S}}(0)=0$, which can be obtained from (16) as $h(0)=-\underaccent{\bar}{\mathcal{S}}(0)$. In (17), $t_{1}$ is the exact time instant from which onward the inner layer sliding manifolds become identically equal to zero. In other words, the sliding mode is enforced on the inner-layer sliding manifold at time $t=t_{1}$. The fundamental concept behind eliminating the reaching phase is to ensure global exact-time convergence for the sliding variables and achieve a complete sliding mode response. This, in consequence, bolsters robustness, as sliding mode controllers typically experience diminished robustness during their reaching phase. The joint cooperative maneuver of the defender-evader team is presented in the next theorem.
Theorem 2 (Maneuver-level Cooperation).
Consider the three-body engagement described using (2), the dynamics of the error variable, (12), and the dual-layer sliding manifold in (14) and (16). If the defender-evader team uses maneuver-level cooperation, and the joint cooperative maneuver of the evader-defender team is designed as
$\mathcal{U}=\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}-\dfrac{2\dot{r}_{\mathrm{ED}}\dot{\lambda}_{\mathrm{ED}}}{r_{\mathrm{ED}}}+\dot{g}(t)-\dot{h}(t)-\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}}),$ (18)
@@
Consider a continuous, radially unbounded Lyapunov function candidate $\mathcal{V}_{1}=\frac{1}{2}\bar{\mathcal{S}}^{2}$. Hereafter, we streamline the notation again by dropping the arguments of variables denoting their time dependency. Upon differentiating $\mathcal{V}_{1}$ with respect to time and using the relations in (14) and (16), one may obtain
$\dot{\mathcal{V}}_{1}=\bar{\mathcal{S}}\dot{\bar{\mathcal{S}}}=\bar{\mathcal{S}}\left(\dot{\underaccent{\bar}{\mathcal{S}}}+\dot{h}\right)=\bar{\mathcal{S}}\left(\ddot{\beta}-\dot{g}+\dot{h}\right),$
which can be further written using the expression in (12) and (13) as
$\dot{\mathcal{V}}_{1}$ $=\bar{\mathcal{S}}\left[\dfrac{2\dot{r}_{\mathrm{ED}}\dot{\lambda}_{\mathrm{ED}}}{r_{\mathrm{ED}}}-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}-\dot{g}+\dot{h}\right$.
$\left.+~\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{ED}}}a_{\mathrm{E}}-\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{ED}}}+\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)a_{\mathrm{D}}\right]$
$\dot{\mathcal{V}}_{1}$ $=\bar{\mathcal{S}}\left[\dfrac{2\dot{r}_{\mathrm{ED}}\dot{\lambda}_{\mathrm{ED}}}{r_{\mathrm{ED}}}-\dfrac{2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{r_{\mathrm{DP}}}+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}-\dot{g}+\dot{h}+\mathcal{U}\right]$ (19)
Upon substituting the expression of $\mathcal{U}$ from (18) into (19), the derivative of the Lyapunov function candidate reduces to
$\dot{\mathcal{V}}_{1}=$ $~\bar{\mathcal{S}}\left[-\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}})+\dfrac{\cos\delta_{\mathrm{PD}}}{r_{\mathrm{DP}}}a_{\mathrm{P}}\right]$
$\leq$ $-\left(\mathcal{K}-\dfrac{1}{r_{\mathrm{DP}}}a_{\mathrm{P}}^{\max}\right)\lvert\bar{\mathcal{S}}\rvert~\leq 0,$ (20)
@@
Note that $\mathcal{U}$ in Theorem 2 is the net control authority that the evader-defender pair has to maintain in order to achieve a successful interception of the pursuer. Next, we endeavor to allocate this net control effort to the evader and the defender. Needless to say, there can be several choices for the allocation of the total control effort between the defender and the evader. However, we focus on a weighted control allocation of the net control effort between the evader and the defender, subjected to the affine constraint (13). Towards that, our aim is to minimize the cost function $\mathcal{C}$, which is defined as
$\mathcal{C}:=\sqrt{\left(\dfrac{a_{\mathrm{E}}}{\Sigma_{\mathrm{E}}}\right)^{2}+\left(\dfrac{a_{\mathrm{D}}}{\Sigma_{\mathrm{D}}}\right)^{2}},$ (25)
where $\Sigma_{\ell}>0$ for $\ell\in\{\mathrm{E},\mathrm{P}\}$ are weights, which are used to allocate the net control effort between the evader and the defender. For a choice $\Sigma_{\mathrm{E}}=\Sigma_{\mathrm{D}}=1$, the minimization of $\mathcal{C}$ is equivalent to instantaneously minimizing the $\mathcal{L}_{2}$ norm of the net control effort. The essence of this effort allocation is presented in the next theorem.
@@
The evader-defender team cooperatively ensures the interception of the pursuer with minimum values of the lateral accelerations as
$a_{\mathrm{E}}$ $=\frac{r_{\mathrm{ED}}r_{\mathrm{DP}}\cos\delta_{\mathrm{ED}}2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}}{\Sigma^{2}r_{\mathrm{DP}}^{2}\cos^{2}\delta_{\mathrm{ED}}+\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)^{2}}$
$-\frac{r_{\mathrm{DP}}^{2}\cos\delta_{\mathrm{ED}}2\dot{r}_{\mathrm{ED}}\dot{\lambda}_{\mathrm{ED}}}{\Sigma^{2}r_{\mathrm{DP}}^{2}\cos^{2}\delta_{\mathrm{ED}}+\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)^{2}}$
$+\frac{r_{\mathrm{ED}}r_{\mathrm{DP}}^{2}\cos\delta_{\mathrm{ED}}\left(\dot{g}(t)-\dot{h}(t)-\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}}\right)}{\Sigma^{2}r_{\mathrm{DP}}^{2}\cos^{2}\delta_{\mathrm{ED}}+\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)^{2}},$ (26)
$a_{\mathrm{D}}$ $=-\dfrac{\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)\left(r_{\mathrm{ED}}2\dot{r}_{\mathrm{DP}}\dot{\lambda}_{\mathrm{DP}}\right)}{\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{ED}}\Sigma\right)^{2}+\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)^{2}}$
$+\dfrac{\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)\left(r_{\mathrm{DP}}2\dot{r}_{\mathrm{ED}}\dot{\lambda}_{\mathrm{ED}}\right)}{\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{ED}}\Sigma\right)^{2}+\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)^{2}}$
$-\dfrac{\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)r_{\mathrm{DP}}r_{\mathrm{ED}}\left(\dot{g}(t)-\dot{h}(t)\right)}{\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{ED}}\Sigma\right)^{2}+\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)^{2}}$
$+\dfrac{\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)\left(r_{\mathrm{ED}}r_{\mathrm{DP}}\mathcal{K}\mathrm{sign}(\bar{\mathcal{S}})\right)}{\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{ED}}\Sigma\right)^{2}+\left(r_{\mathrm{DP}}\cos\delta_{\mathrm{DE}}+r_{\mathrm{ED}}\cos\delta_{\mathrm{DP}}\right)^{2}}.$ (27)
@@
$\mathcal{U}=\mathfrak{B}_{1}a_{\mathrm{E}}-\mathfrak{B}_{2}a_{\mathrm{D}}\implies a_{\mathrm{E}}=\dfrac{\mathcal{U}+\mathfrak{B}_{2}a_{\mathrm{D}}}{\mathfrak{B}_{1}},$ (28)
where $\mathfrak{B}_{1}=\dfrac{\cos\delta_{\mathrm{ED}}}{r_{\mathrm{ED}}}$ and $\mathfrak{B}_{2}=\left(\dfrac{\cos\delta_{\mathrm{DE}}}{r_{\mathrm{ED}}}+\dfrac{\cos\delta_{\mathrm{DP}}}{r_{\mathrm{DP}}}\right)$. Now substituting the value of $a_{\mathrm{E}}$ from (28) in the cost function (25), one may obtain
@@
$\mathcal{C}=\sqrt{\left(\dfrac{\mathcal{U}+\mathfrak{B}_{2}a_{\mathrm{D}}}{\mathfrak{B}_{1}\Sigma_{\mathrm{E}}}\right)^{2}+\left(\dfrac{a_{\mathrm{D}}}{\Sigma_{\mathrm{D}}}\right)^{2}},$ (29)
Computing the partial derivative of $\mathcal{C}$ with respect to $a_{\mathrm{D}}$ and equating it to zero leads to
@@
$\dfrac{\partial\mathcal{C}}{\partial a_{\mathrm{D}}}=\frac{1}{2\mathcal{C}}\left[2\left(\frac{\mathcal{U}+\mathfrak{B}_{2}a_{\mathrm{D}}}{\mathfrak{B}_{1}\Sigma_{\mathrm{E}}}\right)\frac{\mathfrak{B}_{2}}{\mathfrak{B}_{1}\Sigma_{\mathrm{E}}}+2\left(\frac{a_{\mathrm{D}}}{\Sigma_{\mathrm{D}}}\right)\frac{1}{\Sigma_{\mathrm{D}}}\right]=0$
$\implies\frac{\dfrac{\mathfrak{B}_{2}\left(\mathcal{U}+\mathfrak{B}_{2}a_{\mathrm{D}}\right)}{\mathfrak{B}_{1}^{2}\Sigma_{\mathrm{E}}^{2}}+\dfrac{a_{\mathrm{D}}}{\Sigma_{\mathrm{D}}^{2}}}{\sqrt{\left(\dfrac{\mathcal{U}+\mathfrak{B}_{2}a_{\mathrm{D}}}{\mathfrak{B}_{1}\Sigma_{\mathrm{E}}}\right)^{2}+\left(\dfrac{a_{\mathrm{D}}}{\Sigma_{\mathrm{D}}}\right)^{2}}}=0$
$\implies\mathfrak{B}_{2}^{2}\Sigma_{\mathrm{D}}^{2}a_{\mathrm{D}}+\mathfrak{B}_{1}^{2}\Sigma_{\mathrm{E}}^{2}a_{\mathrm{D}}=-\mathfrak{B}_{2}\Sigma_{\mathrm{D}}^{2}\mathcal{U}$
$\implies a_{\mathrm{D}}=\dfrac{-\mathfrak{B}_{2}\Sigma_{\mathrm{D}}^{2}\mathcal{U}}{\mathfrak{B}_{1}^{2}\Sigma_{\mathrm{E}}^{2}+\mathfrak{B}_{2}^{2}\Sigma_{\mathrm{D}}^{2}}=-\dfrac{\mathfrak{B}_{2}\mathcal{U}}{\mathfrak{B}_{1}^{2}\Sigma^{2}+\mathfrak{B}_{2}^{2}},$